OpenStudy (anonymous):

A square tile has an area of 49 square inches. What is the length of this tile?

OpenStudy (domebotnos):

The length and width of a square is the same so the equation to solve would be: 49 = l*w or replace it as 49 = x^2 49 = x^2 √49 = √x^2 7 = x The length and also width is 7. Do u understand? :)

OpenStudy (anonymous):

yes i copmetely undersatnd so the answer is 7?

OpenStudy (domebotnos):

Yes, for the length and width :)

OpenStudy (anonymous):

okay kina says hi